OpenStudy (anonymous):

Can 3x^3+15x^2 be factored?

OpenStudy (anonymous):

3x^2 (x+5)

OpenStudy (anonymous):

Thankyou so much :)

OpenStudy (anonymous):

Yes it can be factored. To factor it, we find the prime factors of \[ 3x^3\] and \[ 15x^2 \] the factors of 3 is 3 *x * x the factor of 15 is 3 * 5 * x * x Now we know the greatest common factor is \[ 3^2 \] Now that we know this we factor it out like below \[ \frac{3x^2 + 15x^2}{3x^2} = 3x^2( 1 + 5)\] Johnny made a mistake and gave you 3x^2(x +5) which is wrong because 3x^2(x+5) = 3^3 + 15x^2

OpenStudy (anonymous):

Okay so it equals 3x^2(1+5), then how would I find the zeros of this equatons. So that I can solve the inequality.